OpenStudy (anonymous):

How do I solve this question 4x - 5y =17 & 3x + 4y = 5 using elimination?

OpenStudy (tkhunny):

The usual ides is rather like a common denominator. We really need a common multiple of something. You can pick the coefficients fo x or those of y. x- coefficients 4 and 3. What is the least common multiple of 4 and 3?

OpenStudy (anonymous):

12?

OpenStudy (tkhunny):

Perfect. The task is to alter the two equations so that the solution is unaltered but something matches up conveniently. Original 4x - 5y =17 <== Multiply this by 3 3x + 4y = 5 <== Multiply this by 4 Altered 12x - 15y =51 12x + 16y = 20 See how that works? We now have the same coefficient for x in both equations.

OpenStudy (anonymous):

Oooh ok I see now ok hold on

OpenStudy (anonymous):

I got a decimal idk y I think I did something wrong

OpenStudy (tkhunny):

Please show your intermediate results. What's wrong with fractions or decimals? What did you get when you subtracted the two altered equations?

OpenStudy (anonymous):

Y=31

OpenStudy (tkhunny):

That's no good. This is why I like to do it a little differently. Subtraction, it turns out, is too complicated. Multiply one of the equations by -1 and then ADD. It is much harder to make errors. This and subtract? No 12x - 15y = 51 12x + 16y = 20 This and ADD?o 12x - 15y = 51 -12x - 16y = -20 Okay, now what do you get when you add the slightly additionally altered equations?

OpenStudy (anonymous):

Is the result gonna be 3, 1?

OpenStudy (anonymous):

-1 sorry

OpenStudy (tkhunny):

Please show your work. Don't leave me guessing what you did. It is so much more difficult to have a helpful conversation. What did you get after ADDING the equations?

OpenStudy (anonymous):

Ok

OpenStudy (tkhunny):

Did the addition thing work better for you AND make sense?

OpenStudy (anonymous):

no and theres something I did wrong

OpenStudy (anonymous):

Hold on tho

OpenStudy (anonymous):

I got 31y = 30

OpenStudy (tkhunny):

You keep dropping that negative sign. Poor thing. Eventually, you will hurt its feelings. :-)

OpenStudy (anonymous):

Thanx for helping